Find the orthogonal and Normal canonical reduction of the quadratic form 7x^2 +6xy+7y^2. Hence identify the conic represented by 7x^2+6xy+7y^2=200.. also find the principal axes of the given quadratic form.
1
Expert's answer
2021-08-17T18:02:12-0400

Orthogonal reduction

F(x,y)=(x,y) "\begin{bmatrix}
 7 & 3 \\
 3 & 7
\end{bmatrix}" "\begin{pmatrix}
 x \\
 y
\end{pmatrix}" ;

"\begin{vmatrix}
 7-k & 3 \\
 3 & 7-k
\end{vmatrix}" =k2-14"\cdot" k+40=0;

k1=4, k2=10;

1)k=4

(3 3)"\cdot\begin{pmatrix}
 x \\
 y
\end{pmatrix}" ="\begin{pmatrix}
 0 \\
 0
\end{pmatrix}"

x+y=0;x=-1,y=1;|(-1 1)}="\sqrt{2}" ;

v1="\frac {1} {\sqrt{2}}\cdot\begin{pmatrix}
 -1 \\
 1
\end{pmatrix}" ;

2) k=10

(-3 3)"\cdot\begin{pmatrix}
 x \\
 y
\end{pmatrix}" ="\begin{pmatrix}
 0 \\
 0
\end{pmatrix}"

-x+y=0;

x=1,y=1;

v1="\frac {1} {\sqrt{2}}\cdot\begin{pmatrix}
 1 \\
 1
\end{pmatrix}" ;

{v1,v2}-new ortogonal basis;

P="\frac {1} {\sqrt{2}}\cdot\begin{pmatrix}
 1 &-1 \\
 1 & 1
\end{pmatrix}"

We make substitution

"\begin{pmatrix}
 \hat{x }\\
 \hat{ y }
\end{pmatrix}" =P-1"\cdot\begin{pmatrix}
 x \\
 y
\end{pmatrix}" ="\frac {1} {\sqrt{2}}\cdot\begin{pmatrix}
 1 &1 \\
 -1 & 1
\end{pmatrix}" "\cdot\begin{pmatrix}
 x \\
 y
\end{pmatrix}" "=\frac {1} {\sqrt{2}}\cdot\begin{pmatrix}
 x+y \\
 -x+y
\end{pmatrix}" ;

q(x,y)=("\hat{x} , \hat{y}" )"\cdot" "P^{T}\cdot\begin{pmatrix}
 7 &3\\
 3 &7 
\end{pmatrix}\cdot P\cdot""\begin{pmatrix}
 \hat{x }\\
 \hat{ y }
\end{pmatrix}"=("\hat{x} , \hat{y}" )"\cdot""\begin{pmatrix}

 10 & 0 \\

 0& 4

\end{pmatrix}" "\cdot" "\begin{pmatrix}
 \hat{x }\\
 \hat{ y }
\end{pmatrix}"=

=10 "\cdot\hat{x}^{2}+4\cdot\hat{y}^2=200" ;

"{\hat{x}^{2}\over \sqrt{20}^{2}}+{\hat{y}^{2}\over \sqrt{50}^{2}}=1" it is ellipse.


Principal axes are :

L1||e1="\frac {1} {\sqrt{2}}\cdot\begin{pmatrix}
 1 \\
 1
\end{pmatrix}" ;

L2||e2="\frac {1} {\sqrt{2}}\cdot\begin{pmatrix}
 -1 \\
 1
\end{pmatrix}" ;

Canonical reducrion:

7x^2+6xy+7y^2=7"\cdot(x^{2}+\frac {6}{7}\cdot x\cdot y+y^{2})="

=7"\cdot(x+\frac{3}{7}\cdot y)^{2}+\frac{40}{7}\cdot y^{2}=200" ;

"\hat{x}=x+\frac{3}{7}\cdot y, \hat{y}=y;"

"7\cdot \hat{x}^2+\frac{40}{7}\cdot \hat{y}^{2}=200;"

"{\hat{x}^2\over\sqrt{\frac {200}{7}}^2}+{\hat{y}^2\over\sqrt{35}^2}=1" - ellipse